### v1

# Next-Embedding Prediction Makes Strong Vision Learners

Abstract
Inspired by the success of generative pretraining in natural language, we ask whether the same principles can yield strong self-supervised visual learners. Instead of training models to output features for downstream use, we train them to generate embeddings to perform predictive tasks directly. This work explores such a shift from learning representations to learning models. Specifically, models learn to predict future patch embeddings conditioned on past ones, using causal masking and stop gradient, which we refer to as Next-Embedding Predictive Autoregression (NEPA). We demonstrate that a simple Transformer pretrained on ImageNet-1k with next embedding prediction as its sole learning objective is effective—no pixel reconstruction, discrete tokens, contrastive loss, or task-specific heads. This formulation retains architectural simplicity and scalability, without requiring additional design complexity. NEPA achieves strong results across tasks, attaining 83.8% and 85.3% top-1 accuracy on ImageNet-1K with ViT-B and ViT-L backbones after fine-tuning, and transferring effectively to semantic segmentation on ADE20K. We believe generative pretraining from embeddings provides a simple, scalable, and potentially modality-agnostic alternative to visual self-supervised learning.
[…]
1 Introduction
Visual pretraining is one of the core topic in computer vision. Self-supervised learning has become the cornerstone of modern visual pretraining method, enabling scalable visual learners without manual labels. At its core, the objective is to learn representations: models are optimized to map raw pixels to a fixed-dimensional representation, which can later be used or fine-tuned for downstream tasks. This philosophy unites methods based on instance discrimination [chen2020simclr, he2020moco], self-distillation [caron2021dino], and masked reconstruction [he2022mae, bao2022beit]. The goal is to learn visual representations that can be consumed by downstream modules at various scales, from lightweight task-specific heads to large cascaded systems such as vision-language models.
Figure 1: Next-Embedding Predictive Autoregression (NEPA). An image is split into patches and embedded into a sequence. An autoregressive model predicts the next embedding from previous ones, mirroring next-token prediction in language models.
A fundamentally different paradigm has underpinned the success of modern natural language processing. Language models are not pretrained to be feature extractors; they are trained to be generative and predictive systems. The objective is not to produce a static embedding of a sentence, but to model the data distribution itself through a simple, causal objective [radford2018gpt, radford2019gpt2, brown2020gpt3, achiam2023gpt4]. This training compels the model to internalize the semantics and conditional dependencies within language. Inference is no longer a two-stage process of encoding followed by task-solving, but a single predictive computation carried out by the model itself. This distinction is fundamental. It suggests that generative prediction, rather than representation learning, may offer a direct route to scale up pretraining. A very recent line of research has shifted to this philosophy, e.g., early pixel-level generative pretraining (iGPT) showed transferable features but struggled with very long sequences and weak semantic alignment [chen2020igpt]. JEPA [assran2023jepa] moves beyond pixels by predicting latent targets and aligning more closely with semantic structure. Still, JEPA trains by regressing to latent targets from a momentum encoder, rather than using generative prediction as the self-supervised objective.
We are thus tempted to ask whether minimal causal pretraining might also give rise to strong vision learners. Concretely, an image is decomposed into patches, which are mapped into a sequence of patch-wise embeddings. A causal transformer is then trained to predict the next embedding given all previous ones, closely mirroring the next-token prediction paradigm in language models. We use a stop-gradient on the target embeddings to create a stable predictive task. This formulation is deliberately minimalist. It requires no pixel-level decoder, no discrete visual tokenizer, and none of the engineered data augmentations, negative pairs, or momentum encoders common in contrastive learning. The entire learning signal is derived from the model’s ability to anticipate the future in embedding space. We call this new family of models Next-Embedding Predictive Autoregression (NEPA).
The effectiveness of this simplicity is demonstrated through its performance. A standard Vision Transformer, pretrained on ImageNet-1K using only our next-embedding prediction objective, achieves state-of-the-art classification accuracy after standard fine-tuning. Furthermore, the learned model transfers effectively to dense prediction tasks, achieving strong results on ADE20K semantic segmentation. These results validate that a purely predictive model, free from the architectural complexity of other paradigms, can learn the rich, contextual features required for diverse visual tasks.
2 Related Work
Visual Self-Supervised Learning.
Self-supervised learning has become the leading paradigm for visual learning, leveraging large unlabeled datasets to learn representations without annotations [wang2015video, noroozi2016puzzles, zhang2016color, pathak2017move, gidaris2018rotation]. Early progress was driven by contrastive and self-distillation approaches [hadsell2006dimensionality, wu2018unsupervised, oord2018cpc, hjelm2018dim, bromley1993siamese, he2020moco, chen2021mocov3, chen2020simsiam, chen2020simclr, grill2020byol, caron2021dino], which learn invariances across augmented views but often require large batches or memory banks. A parallel line uses reconstruction objectives, where lightweight decoders recover masked pixels or tokens at scale [vincent2008dae, pathak2016context, devlin019bert, he2022mae, bao2022beit]. The limitations of reconstruction have motivated predictive representation learning, which forecasts semantic embeddings rather than raw inputs, as in JEPA [assran2023jepa]. Yet JEPA remains fundamentally representation-centric: its pretrained encoder produces features that downstream modules must consumed with separate heads or fine-tuning, rather than directly driving task behavior through prediction itself.
Generative Pretraining.
Generative pretraining has driven major advances in language [radford2018gpt, radford2019gpt2, brown2020gpt3, achiam2023gpt4] and has a long visual lineage. Early pixel-level generative pretraining (iGPT) showed transferable features but struggled with very long sequences and weak semantic alignment [chen2020igpt]. Subsequent work introduced discrete tokenizers [van2017vqvae, esser2021vqgan, lee2022rqvae, sun2024llamagen, pang2024randar] to compress images into short code sequences amenable to transformer generative pretraining [esser2021vqgan, ramesh2021dalle, lee2022rqvae], enabling large-scale image generation. Recent scaling efforts, such as VAR [tian2024var] and LlamaGen [sun2024llamagen], further close or surpass the diffusion modeling performance. In parallel, tokenizer-free generative pretraining operates directly in continuous spaces [fan2024fluid, li2024mar], and masked autoregression blends bidirectional context with ordered generation for efficiency and quality [li2024mar, wu2025dcar, chang2022maskgit, li2023mage]. Unlike these works, we treat prediction itself as the learning signal and operate directly in the embedding space without a generative decoder.
Representation Prediction.
Predicting representations instead of raw signals is a long-standing idea in both neuroscience and machine learning. Predictive coding suggests that the brain forecasts latent causes of sensory inputs [rao1999predictive], while Contrastive Predictive Coding (CPC) and related frameworks apply this principle in practice [oord2018cpc, wu2018unsupervised, hjelm2018dim, bromley1993siamese, he2020moco, chen2021mocov3, chen2020simsiam, chen2020simclr, grill2020byol, caron2021dino]. JEPA makes this explicit by predicting target embeddings from a context view in a shared space [assran2023jepa], and CLIP can be interpreted as cross-modal embedding prediction aligning text and image representations [radford2021clip]. However, current JEPA-style systems typically rely on auxiliary encoders and prediction heads and predict in parallel rather than causally, contrasting with the simple, scalable next-token objective that underpins generative pretrained language models, which is a formulation proven both effective and efficient.
3 Method
In this section, we first introduce the next-embedding prediction objective, followed by the architectural design that enables both pretraining and downstream fine-tuning.
3.1 Next Embedding Prediction
Given an input image $x$, we divide it into $T$ non-overlapping patches and map each patch to an embedding using a shared encoder $f$, resulting in a sequence of embeddings $z=\{z_{1},z_{2},\ldots,z_{T}\}$. We consider next embedding prediction as the training objective, where an autoregressive predictor $h_{\theta}$ aims to model the next embedding conditioned on the previous ones:
$\widehat{z}_{t+1}=h_{\theta}(z_{\leq t}),$ (1)
This is directly analogous to next-token prediction in language modeling, but operates entirely in continuous embedding space rather than over discrete tokens or pixel values.
To optimize this objective, we adopt a similarity-based loss inspired by SimSiam [chen2020simsiam]. At each position $t$, the model predicts the next embedding $\widehat{z}_{t+1}$ based on the context $z_{\leq t}$, and compares it to the corresponding embedding $z_{t+1}$ obtained directly from the encoder. We treat $\widehat{z}_{t+1}$ as the query and $z_{t+1}$ as the target. Both vectors are normalized to unit length, and the similarity is measured via negative cosine similarity:
$\mathcal{D}(z,\widehat{z})=-\frac{1}{T-1}\sum_{t=1}^{T-1}(\frac{z_{t+1}}{\|z_{t+1}\|_{2}}\cdot\frac{\widehat{z}_{t+1}}{\|\widehat{z}_{t+1}\|_{2}}).$ (2)
To avoid degenerate solutions, we follow chen2020simsiam and apply a stop-gradient operation to the target embeddings. The final training loss is defined as:
[…]
Intuitively, this objective encourages the model to predict embeddings that are semantically aligned with the ground truth, without requiring explicit reconstruction of the input. The training can be easily implemented using Algorithm 1.
Algorithm 1 Next-Embedding Prediction
# f: embedding layer
# h: autoregressive model
for pixel_values in loader: # x, [B, H, W, C]
input_embed = f(pixel_values) # z, [B, T, D]
pred_embed = h(input_embed) # z_hat, [B, T, D]
loss = D(input_embed, pred_embed) # loss
loss.backward() # back-propagate
update(f.param, h.param) # update parameters
def D(z, z_hat):
target = z.detach() # stop gradient
pred = z_hat[:, 0:T-1, :] # shift, [B, T-1, D]
target = target[:, 1:T, :] # shift, [B, T-1, D]
# Use any suitable distance metric.
pred = normalize(pred, axis=-1) # l2-norm
target = normalize(target, axis=-1) # l2-norm
return -(pred * target).sum(dim=-1).mean()
3.2 Model Architecture
We adopt a standard Vision Transformer (ViT) backbone [dosovitskiy2020vit] with causal attention masking. Unlike pixel-level reconstruction methods [he2022mae, bao2022beit], our approach requires no separate decoder. The Transformer predicts future patch embeddings directly from past ones, using a single backbone for both context encoding and prediction, similar to autoregressive language models. Images are split into non-overlapping patches using a Conv2d patch embedding layer, with learnable positional embeddings added before being fed into the Transformer. We adopt a pre-norm design with LayerNorm [ba2016layernormalization] and apply a final LayerNorm to the output features. To improve stability and scalability, we incorporate modern training and normalization practices inspired by DINOv3 [siméoni2025dinov3] and VisionLLaMA [chu2024visionllama], as shown in Figure 2. These modeling designs are helpful for training but orthogonal to our core framework, and are included below for reproducibility and completeness.
RoPE.
We adopt Rotary Position Embedding (RoPE [su2024roformer]) at all layers to encode relative positions via complex rotations in attention. RoPE improves generalization and positional reasoning over varying sequence lengths [touvron2023llama, bai2023qwentechnicalreport, siméoni2025dinov3].
LayerScale.
We adopt LayerScale [touvron2021cait] to stabilize training by applying learnable per-channel scales (initialized to $10^{-5}$) to residual branches, thereby improving convergence with minimal computational overhead.
SwiGLU.
We replace the standard GeLU activation [hendrycks2017gelu] in vision transformer feed-forward networks [dosovitskiy2020vit] with the SwiGLU activation [shazeer2020glu]. While our experiments (Section 4) show only modest improvements over GeLU, we retain SwiGLU to align with recent architectures [oquab2023dinov2, siméoni2025dinov3, chu2024visionllama, sun2024llamagen] and to ensure compatibility with prevailing designs in both advanced vision models and large language models [touvron2023llama, bai2023qwentechnicalreport].
QK-Norm.
To further enhance training stability, we adopt query-key normalization (QK-Norm [henry2020qknorm]). This helps mitigate issues such as gradient explosion or collapse in attention and facilitates optimization in deeper transformers [dehghani2023vit22b, esser2024sd3]. In our implementation, we apply LayerNorm [ba2016layernormalization] without bias and without learnable parameters, ensuring a lightweight yet effective normalization scheme.
Figure 2: Images are tokenized via a Conv2d patch embedder before entering a pre-norm Transformer with LayerNorm. Modern stabilization components (RoPE [su2024roformer], LayerScale [touvron2021cait], SwiGLU [shazeer2020glu], and QK-Norm [henry2020qknorm]) are applied at all layers.
3.3 Task-Specific Heads
Vision tasks such as image classification, semantic segmentation, object detection, and generative modeling differ in output structure and supervision signals. Among visual tasks, image classification and semantic segmentation are the most commonly used benchmarks for evaluating pretrained models, as they reflect global and dense prediction capabilities. Thus, in our evaluation, we focus primarily on these two tasks. To evaluate the downstream utility of our pretrained model, we design task-specific heads for classification and segmentation, respectively. These heads are lightweight, compatible with standard practices, and do not modify the model architecture.
Classification Head.
For image-level recognition tasks, we adopt a standard linear classification head [dosovitskiy2020vit, bao2022beit, he2022mae, caron2021dino, assran2023jepa]. Specifically, given the final output hidden states $\{z_{1},z_{2},\ldots,z_{T}\}$ from the pretrained transformer, we select the last embedding $z_{T}$ as a compact representation and apply a linear classifier to predict the category label. We train this head using the cross-entropy loss over the predicted logits $\widehat{y}$ and the ground-truth label $y$.
Segmentation Head.
Following previous work [he2022mae, bao2022beit], for semantic segmentation, we adopt a UPerNet head [xiao2018upernet] that operates on multi-scale features extracted from intermediate transformer blocks. We supervise predictions using the standard pixel-wise cross-entropy loss. To enable spatial context modeling, we disable causal masking and allow bidirectional attention during fine-tuning.
4 Experiments
shifting causal masking stop-grad 50k acc (%)
fail
73.6
fail
[…]
(a) Effects of causal masking, autoregressive shifting, and stop-gradient. Removing shifting leads to divergence during fine-tuning, while removing stop-gradient results in training collapse during pretraining.
LayerScale RoPE QK-Norm GatedMLP 100k acc (%)
78.2
77.4
80.2
fail
81.1
[…]
(b) Ablation of architectural components on 100k-step accuracy.
masking ratio 100k acc (%)
[…]
40 76.4
60 75.7
(c) Effect of input embedding masking ratio on 100k-step pretraining accuracy using a baseline ViT backbone with embedding prediction.
freeze embedding 100k acc (%)
73.1
[…]
(d) Effect of freezing the patch embedding layer in the base model with LayerScale on 100k-step pretraining accuracy.
attn type 100k acc (%)
[…]
causal 81.3
(e) Effect of attention type during fine-tuning for our full model. We compare causal and bidirectional attention on 100k-step pretraining accuracy.
Table 1: Ablations studies. Default settings are marked in gray. Experiments that fail to converge are marked with fail.
We conduct comprehensive experiments to evaluate NEPA of different sizes and configurations. All models are pretrained from scratch on the ImageNet-1K dataset [russakovsky2015imagenet1k] without labels. We begin with a controlled study on the contributions of key components in our next embedding prediction algorithm, including causal masking, temporal shifting, and stop-gradient. We then ablate a series of architectural design choices that further improve performance. In comparison to previous work, we evaluate NEPA on two standard benchmarks: ImageNet-1K [russakovsky2015imagenet1k] for image classification and ADE20K [zhou2017ade20k] for semantic segmentation. To better understand the model’s internal behavior, we examine the learned attention patterns and embedding.
4.1 Ablation Study on Core Algorithm Design
This section separately studies the impact of masking, autoregressive shifting, and stop-gradient, which are core algorithm designs for the next embedding prediction objective. All models are pretrained with a batch size of 2048 for either 50k or 100k steps, and then fine-tuned for downstream tasks; full training details are provided in Appendix D. All reported results are obtained using the exponential moving average (EMA [tarvainen2017mean]) model with an EMA decay rate of $0.9999$. We report the top-1 accuracy on the ImageNet-1K validation set to compare different settings. Tables 1 and 1 outline the results.
Shifting.
In our default setup, the model predicts the embedding of the next token rather than copying the current input. Removing this shift reduces the objective to an identity mapping, which provides no meaningful prediction target. After 50k steps of pretraining, this variant diverges during fine-tuning and fails to converge to a usable model.
Causal masking.
By default, each token attends only to its predecessors, ensuring a causal prediction setup. Removing the causal mask allows every token to attend bidirectionally, effectively turning the model into a reconstruction model. After 50k steps of pretraining, this variant achieves only 73.6% top-1 accuracy after fine-tuning, compared to 76.8% with causal masking.
Stop-gradient.
Following SimSiam [chen2020simsiam], we stop gradients on the ground-truth embeddings to prevent collapse. Without this operation, the training loss collapses to 1-1 during pretraining, indicating all embeddings are the same.
Random masking.
We also applied random masking to the input embeddings, while still predicting all target embeddings. In Table 1, we tested different masking ratios at 100k pretraining steps and observed a clear performance drop as masking increased: 0% masking achieved 78.2% top-1 accuracy, 40% masking achieved 76.4%, and 60% masking achieved 75.7%. This result highlights a key difference from masked image modeling paradigms such as MAE, where masking is essential to prevent trivial reconstruction. In contrast, our autoregressive setup naturally avoids shortcut solutions. This suggests that random masking, while useful for preventing trivial reconstruction in pixel-space objectives, is less compatible with embedding-level prediction, where causal modeling already imposes a meaningful learning signal. Masking in this context introduces input corruption, disrupts sequence structure, and creates a training-inference mismatch—none of which are necessary when the prediction task is inherently non-trivial. These findings reinforce the appeal of our approach: prediction is sufficient, and masking, rather than helping, may obscure the signal.
4.2 Ablation Study on Orthogonal Components
We ablate 4 architectural components (LayerScale, RoPE, QK-Norm, and SwiGLU) introduced in Section 3.2. Starting from a plain ViT backbone, we incrementally add each component and report top-1 accuracy after 100k steps of pretraining followed by fine-tuning. As shown in Table 1, each component contributes to performance gains, and their combination yields the best result. Notably, enabling RoPE significantly boosts accuracy, while QK-Norm further improves stability and final performance.
LayerScale.
We employ LayerScale to enhance training stability during pretraining. Although LayerScale slightly reduces top-1 accuracy on ImageNet-1K after fine-tuning, we find it plays a crucial role in stabilizing optimization and helping the model converge faster to a lower training loss during pretraining (Figure 3). We attribute the degradation in fine-tuning performance to LayerScale’s tendency to slow convergence in the supervised stage due to reduced gradient magnitudes. To address this, we freeze the patch embedding layer during fine-tuning, which accelerates convergence and improves overall performance, in line with previous observations [touvron2021cait, shazeer2020glu]. For these reasons, we retain LayerScale in our final model. We note that this trade-off between pretraining stability and fine-tuning speed is a common theme in modern architectures.
RoPE.
We use RoPE [su2024roformer] to encode positional information in a continuous and extrapolatable manner. Compared to absolute position embeddings, RoPE leads to a significant improvement in top-1 accuracy after fine-tuning.
QK-Norm.
We use QK-Norm [henry2020qknorm] to stabilize attention and prevent gradient explosion during pretraining with SwiGLU as Section 3.2. Without QK-Norm, we observe that the model becomes unstable when combined with SwiGLU, either failing to converge or requiring a significantly smaller learning rate to avoid divergence. However, reducing the learning rate leads to underfitting and degraded performance. By applying QK-Norm to the query and key projections, we ensure smooth optimization even in deeper configurations, allowing the model to benefit from SwiGLU without sacrificing training stability.
SwiGLU.
We adopt the Gated MLP variant with SwiGLU activation [shazeer2020glu], following the design used in many recent vision [oquab2023dinov2, siméoni2025dinov3, chu2024visionllama, sun2024llamagen] and language models [touvron2023llama, bai2023qwentechnicalreport]. We find that replacing GeLU with SwiGLU yields only marginal performance improvements. Despite its limited empirical gain in our setting, we retain SwiGLU to maintain architectural alignment with modern transformer-based backbones in both vision and language domains. This choice ensures better compatibility with emerging model designs and facilitates future integration into unified frameworks.
Figure 3: Ablation of key components in NEPA pretraining. Left: EMA accuracy with and without AR shift. Without the autoregressive shift, training diverges early. Middle-left: Training loss with and without stop-grad; removing stop-grad causes representation collapse. Middle-right: Training loss with and without LayerScale; LayerScale stabilizes optimization and accelerates convergence. Right: Gradient norm with and without QK-Norm; QK-Norm suppresses gradient explosion and improves smoothness.
4.3 Comparisons with Previous Results
In this section, we evaluate pretrained NEPA checkpoints on two standard vision tasks: single-label image classification and semantic segmentation. For classification, we fine-tune on ImageNet-1K [russakovsky2015imagenet1k] and report top-1 accuracy. For segmentation, we attach a UperNet [xiao2018upernet] decoder to the NEPA backbone and fine-tune on ADE20K [zhou2017ade20k], reporting mean Intersection-over-Union (mIoU). We experiment with both Base and Large model variants, each pretrained on ImageNet-1K with a patch size of 14 (following DINOv2 [oquab2023dinov2]) and a global batch size of 4096 (following MAE [he2022mae]). We use the same pretrained checkpoint for all downstream evaluations. Models are trained using the HuggingFace Trainer with distributed data parallelism (DDP) on 8 NVIDIA H100 GPUs. The Base model is trained for 1600 epochs (approximately 3 days), and the Large model for 800 epochs (approximately 5 days). Full training and fine-tuning details are provided in Appendix D.
Scaling behavior.
We show that our next embedding prediction framework scales effectively with the size of the model. As shown in Figure 4, model accuracy steadily improves with increased training steps, and we observe no signs of overfitting even under extended compute budgets. Beyond verifying pretraining stability, the scaled models serve as the foundation for subsequent analysis of attention behavior and embedding structure, as well as for evaluating downstream performance via standard fine-tuning on classification and segmentation tasks.
Fine-tuning practice.
We find that enabling bidirectional attention during fine-tuning improves classification performance on ImageNet-1K when the model is pretrained for 100k steps, as shown in Table 1. Nonetheless, NEPA achieves competitive results even under causal attention, indicating that autoregressive embeddings retain sufficient global information for downstream tasks. To remain consistent with the autoregressive formulation, we use causal attention as the default setting in our experiments. However, we also report the results of bidirectional attention for completeness. For semantic segmentation, we adopt bidirectional attention during fine-tuning by default, as each output embedding corresponds to a localized prediction and requires access to the full spatial context of the input image.
Classification results.
We follow standard protocols [touvron2021deit, bao2022beit, he2022mae] and fine-tune NEPA on the labeled ImageNet-1K dataset. We use similar hyperparameters as MAE [he2022mae], including layer-wise learning rate decay [clark2020llrd, bao2022beit], RandAugment [dubuk2020randaugement], label smoothing [szegedy2016label], mixup [zhang2018mixup], cutmix [yun2019cutmix], and DropPath [huang2016droppath]. Results are summarized in Table 2. Compared to prior pretraining methods such as MAE [he2022mae], BEiT [bao2022beit], and MoCo [he2020moco], NEPA achieves competitive top-1 accuracy, reaching 83.8% on the Base model and 85.3% on the Large model. Our method does not require any task-specific head, decoder, or auxiliary loss. Unlike contrastive or masked prediction approaches, we adopt a single-stream autoregressive formulation with one forward pass, no reconstruction target, and no multi-branch architecture. Moreover, the strong performance under causal attention demonstrates the effectiveness of embedding-level autoregression in capturing transferable semantics.
Figure 4: ImageNet-1K validation Top-1 accuracy versus training epochs. For each epoch’s checkpoint, we perform a lightweight hyperparameter search and report the best accuracy. Fine-tuning uses causal attention. The top plot corresponds to the base model, and the bottom plot to the large model.
Model Pretrain Task Pretrain Framework Decoder # FWD / step Epochs Acc (%)
ViT-B
MoCo v3-B [chen2021mocov3] contrastive learning siamese mlp proj. head 2 600 83.2
BEiT-B [bao2022beit] masked token pred masked modeling linear pred. head 1 800 83.4
DINO-B [caron2021dino] self-distillation siamese mlp proj. head N 1600 83.6
MAE-B [he2022mae] masked pixel pred masked autoencoder transformer decoder 1 1600 83.6
[…]
ViT-L
MoCo v3-L [chen2021mocov3] contrastive learning siamese mlp proj. head 2 600 84.1
iBot-L [zhou2022ibot] self-dist & masked token pred siamese & masked modeling mlp proj. head 4 1000 84.8
BEiT-L [bao2022beit] masked token pred masked modeling linear pred. head 1 800 85.2
MAE-L [he2022mae] masked pixel pred masked autoencoder transformer decoder 1 1600 85.6
JEPA-L [assran2023jepa] masked embed pred siamese & masked modeling transformer predictor 2 300 85.2
[…]
Table 2: Comparison of different self-supervised learning frameworks on ImageNet-1K classification. Results are grouped by model scale, with Base models in the upper block and Large models in the lower block. Effective pretraining epochs are used based on the actual number of images or views seen during training; see [zhou2022ibot] for details. indicates methods that use causal attention during fine-tuning. Denotes results based on our implementation.
Method Pre-train data ViT-B ViT-L
Supervised IN1K w/ labels 47.4 49.9
MoCo v3 [chen2021mocov3] IN1K 47.3 49.1
BEiT [bao2022beit] IN1K + DALLE 47.1 53.3
MAE [he2022mae] IN1K 48.1 53.6
[…]
Table 3: Comparison of ADE20K semantic segmentation (mIoU) under different pretraining methods.
Semantic Segmentation.
We evaluate NEPA on ADE20K by attaching a standard UperNet [xiao2018upernet] decoder and fine-tuning the model using bidirectional attention, as described before. ADE20K is a challenging scene parsing benchmark with dense pixel-wise annotations. We follow the training recipe from MMSegmentation [mmseg2020], using a crop size of $512\times 512$, batch size of 16, and 160K training iterations. As shown in Table 3, NEPA achieves competitive performance with only ImageNet-1k pretraining, reaching 48.3% mIoU for Base model and 54.0% mIoU for Large model. Despite not using any decoder or pixel-level objective during pretraining, NEPA transfers effectively to dense prediction tasks. Unlike prior work relying on contrastive learning or pixel reconstruction, our method learns transferable representations purely through next embedding prediction.
5 Quantitative Results
In this section, we investigate how NEPA organizes visual information by analyzing its attention patterns and the structure of its learned embeddings. Our goal is to understand whether the next-embedding prediction objective induces meaningful global dependencies and semantic organization.
5.1 Attention Map Analysis
To better understand how NEPA exploits context when predicting future patches, we visualize representative attention maps on ImageNet-1K images in Figure 5(a). For each triplet, the first column marks the current query patch in the original image, and the second column shows the corresponding attention maps from NEPA.
Interestingly, the attention maps are often long-ranged and object-centric. Across a wide range of categories and viewpoints, the model consistently allocates most of its attention to regions that are semantically related to the query patch. These regions typically include other parts of the same object or nearby informative structures, rather than uniformly attending to all patches or focusing only on local neighbors. For example, when the query patch lies on the head of an animal or the body of a person, the model tends to attend to other body parts and task-relevant background regions, even when they are spatially distant. In cluttered scenes, NEPA suppresses distractors and concentrates on a small subset of visually coherent patches. These qualitative results suggest that the predictive objective encourages the Transformer to form global, semantically meaningful dependencies between patches, effectively learning to localize and group object parts without any explicit supervision.
5.2 Embedding Analysis
We further analyze the embeddings produced by NEPA by examining the similarity between the predicted embedding of the next patch and all other patch embeddings within the same image, shown in the third column of each triplet in Figure 5(a). We observe that the predicted embedding is most similar to patches that belong to the same object or semantic region as the current patch, while unrelated background areas exhibit much lower similarity. This behavior emerges even though NEPA is never trained with explicit labels or region annotations.
(a) ImageNet-1K validation samples (unseen during pretraining).
(b) MSCOCO validation samples (out of distribution during pretraining).
Figure 5: Attention and embedding analyses. Each example consists of three views: (i) the query patch ($\square$) highlighted in the original image, (ii) the attention map from the NEPA showing which patches the model attends to when predicting the next embedding, and (iii) the embedding-similarity map showing the cosine similarity between the predicted embedding and all other patch embeddings in the same image. Warmer colors indicate higher attention or greater similarity; cooler colors indicate lower values.
The embedding similarity maps reveal that the model does not merely memorize local texture; instead, it learns to extrapolate object-level structure. When the query patch lies on a distinctive part of an object, the predicted embedding is highly similar to distant patches covering other parts of that object, and often to patches that are occluded or outside the local neighborhood. In homogeneous backgrounds, the similarity patterns become more diffuse but remain coherent, reflecting uncertainty in the visual context. Overall, these observations indicate that generative pretraining with next-embedding prediction encourages NEPA to learn embeddings that are both predictive and semantically organized, which likely contributes to its strong downstream transfer performance.
6 Conclusion and Future Work
This work revisits the core idea of causal next-token prediction, not in pixel or token space, but in the embedding space of vision models. We demonstrate that simple next embedding prediction is sufficient to learn transferable visual representations that scale. By treating patch embeddings as prediction targets, we avoid brittle design choices tied to handcrafted pretext tasks, and instead rely on the structure implicitly learned through the sequence. With only IN1K self-supervised pretraining, NEPA matches or slightly outperforms mainstream methods on classification and segmentation transfer, while using a simpler training paradigm (single forward pass, no decoder). What emerges is not just an algorithm, but a different perspective: that the simplicity of autoregression, when properly adapted to vision, can help unify pretraining paradigms across modalities.
Modality-agnostic potentials.
Recent Large Language Models [touvron2023llama] increasingly adopt tied embeddings, where the input and output embedding matrices are shared. This mechanism is effectively equivalent to predicting the next embedding in the latent space, which is precisely the principle underlying our framework. From this perspective, our work does not introduce a new paradigm, but rather reveals a unifying view: different modalities can be trained under the same objective, with embeddings serving as a common representational currency.
Generative potentials.
At the same time, our formulation naturally extends toward generative modeling. By coupling NEPA with a suitable image decoder or diffusion-based generator [goodfellow2020gan, ho2020ddpm, song2021ddim, peebles2023dit], the same autoregressive embedding predictor could be used for image synthesis or editing. Exploring this direction, which bridges representation learning and generation within a unified model, remains an exciting avenue for future work.
Acknowledgment
This work is supported in part by a departmental fellowship to Sihan Xu and Weinberg Cognitive Science Fellowship to Sihan Xu and Ziqiao Ma. We acknowledge the computing resources of Google Cloud Platform and Lambda Labs. The authors would like to thank Jiayi Pan and Freda Shi for their valuable feedback.
[…]
Supplementary Material
Appendix A Broader Impacts
NEPA is a generic self-supervised visual pretraining method and, like other large models, may inherit dataset biases or be misused to create misleading visual content, so care is needed in data selection, and deployment, including appropriate safeguards and monitoring.
Appendix B Methodology Comparisons
Relation to CPC [oord2018cpc].
CPC-style methods mainly train an encoder with a contrastive loss, where a small autoregressive module just aggregates context. In NEPA, the main learned object is the transformer predictor itself: we use a simple shared embedding layer and directly regress the next embedding without negatives or a contrastive head, which makes the architecture easier to scale.
Relation to GPT [radford2018gpt, chen2020igpt].
NEPA shares the causal prediction idea of GPT, but works fully in the continuous embedding space. We do not use task-specific decoders, tokenizers, or language heads during pretraining.
Relation to JEPA [assran2023jepa].
JEPA uses separate encoders for context and target views plus a heavy head to score representation pairs. As shown in Fig. 6, NEPA keeps the JEPA-style latent prediction goal but simplifies the architecture to a single embedding layer and an autoregressive transformer predictor, without asymmetric branches or an extra head.
Figure 6: Comparison between JEPA and NEPA.
Appendix C Implementation Details
We first prototyped and validated NEPA using the timm [rw2019timm] library. For large-scale experiments, we switched to the Hugging Face ecosystem, relying on transformers [wolf-etal-2020-transformers], datasets [lhoest-etal-2021-datasets], and evaluate. Our implementation is mainly based on the ViT code in transformers.models.vit.modeling_vit, and the training scripts are adapted from the official run_mim and run_image_classification examples. For data augmentation, we use the timm utilities, in particular Mixup and create_transform. All ImageNet experiments use the ILSVRC/imagenet-1k dataset accessed through Hugging Face datasets. ADE20K experiments follow the standard recipes and data processing pipelines from mmsegmentation [mmseg2020]. Our code and model is available at https://sihanxu.github.io/nepa.
Appendix D Training Details
Pre-training.
The full pre-training hyperparameters are listed in Table 4. For all models we use the standard linear learning-rate scaling rule, i.e., $\text{lr}=\text{base\_lr}\times B/256$, where $B$ is the global batch size.
config value
optimizer AdamW [loshchilov2018adamw]
base learning rate 3e-4
weight decay 0.05
optimizer momentum $\beta_{1},\beta_{2}{=}0.9,0.95$ [chen2020igpt]
batch size 4096
learning rate schedule cosine decay
warmup epochs [goyal2018warmup] 40
augmentation RandomResizedCrop
Table 4: Pre-training setting.
End-to-end Fine-tuning.
Fine-tuning settings are summarized in Table 5. We use the same linear learning-rate scaling rule as in pre-training and apply layer-wise learning-rate decay. Instead of using a fixed decay factor, we linearly increase the decay rate from a smaller value to $1.0$ over the course of training, which helps reduce overfitting at the beginning while still allowing all layers to adapt later.
config value
optimizer AdamW
base learning rate 1e-3
weight decay 0.05
optimizer momentum $\beta_{1},\beta_{2}{=}0.9,0.999$
layer-wise lr decay [bao2022beit, he2022mae] $0.35$ (B) $0.60$ (L)
batch size 1024
learning rate schedule cosine decay
warmup epochs 5
training epochs 100 (B), 50 (L)
augmentation RandAug (9, 0.5) [touvron2021deit, bao2022beit, he2022mae]
label smoothing [szegedy2016label] 0.1
mixup [zhang2018mixup] 0.8
cutmix [yun2019cutmix] 1.0
drop path [huang2016droppath] 0.1 (B) 0.2 (L)
Table 5: Fine-tuning setting.
Appendix E Additional Experiments
E.1 Training Dynamics
We report the pre-training loss curves and the attention maps from NEPA-L of intermediate checkpoints in this section (Figure 7). The plots and figures illustrate the overall training stability of NEPA and how the attention evolves as pre-training progresses.
Figure 7: Pre-training dynamics of NEPA. Top: pre-training loss curves; Bottom: evolution of NEPA-L attention maps.
E.2 Limitations and Failure Cases
Linear Probing.
We evaluated two variants for linear probing, one using the last embedding of the autoregressive model and one using the average of all embeddings, with results reported in Table 6. NEPA performs poorly under standard linear probing. This is expected, as the output representation is very close to the features right after the embedding layer, thus the probing result mainly reflects this shallow representation rather than the full capacity of the predictor.
NEPA-B Last Embed Avg. Embed
Acc (%) 11.3 14.1
Table 6: Linear probing on NEPA-B. Top-1 accuracy using the last autoregressive embedding or the average over all embeddings.
Quantitative Failure Examples.
We also inspect typical failure cases in Figure 8. Under the current scale, NEPA often struggles with images that require non-trivial and reasoning-intensive physical understanding, such as interpreting reflections, shading, and shadows, as well as scenes containing many small or overlapping objects. In such cases, the model tends to produce uncertain or inconsistent predictions, suggesting room for improvement in reasoning about complex spatial layouts. We hypothesize that this reflects the limitation of the ImageNet dataset on which we train our current model [taesiri2023imagenet, neuhaus2023spurious], and this could be addressed as we scale up with more diverse datasets.
Figure 8: Quantitative Failure Examples. In a multi-object scene with strong reflections, the model confuses highlights on metal surfaces with object embeddings; under backlighting, shadowed regions are misinterpreted as trees; for the animal, shaded skin is treated as background; and in the last example, bright reflective regions of the object are also mistaken for background.
E.3 Ablations on MAE Baseline Components
Table 7 reports the effects of adding our orthogonal components to a MAE baseline, where they bring little or no gain in our experiments. We hypothesize masked reconstruction is less sensitive to embedding/positional issues, RoPE mainly helps autoregressive models, and QK-Norm is only useful when paired with LayerScale.
MAE Our impl w/ LayerScale w/ RoPE w/ QK-Norm
Acc (%) 85.6 85.5 85.6 85.6
Table 7: Ablations on MAE with our components. Applying LayerScale, RoPE, and QK-Norm to a strong MAE baseline brings little additional gain on performance.
E.4 Layer-wise LR decay and Overfitting
We study the effect of our layer-wise learning-rate decay (LLRD) schedule on overfitting. Instead of using a fixed decay factor, we linearly increase the decay from a small initial value to $1.0$ over the course of fine-tuning, effectively unfreezing deeper layers gradually. As shown in Table 8, this schedule reduces overfitting and improves performance: on the NEPA-B backbone the accuracy increases from $83.0\%$ to $83.8\%$.
Layer-wise LR decay $0.65$ $0.35\rightarrow 1.00$
Acc (%) 83.0 83.8
Table 8: Effect of layer-wise LR decay on NEPA-B. Comparing a fixed decay factor of $0.65$ with a schedule that increases from $0.35$ to $1.00$ during fine-tuning.
Figure 9: Additional attention and embedding visualizations.
E.5 Reproduced Results without SwiGLU
Finally, we evaluate NEPA without SwiGLU, replacing it with standard feed-forward layers. For this comparison we lightly tune the learning rate and layer-wise LR decay and report the best setting. Table 9 shows that we can reproduce similar results in this configuration, indicating that SwiGLU is non-essential optimization and that the main gains come from the overall predictive architecture.
NEPA-B w/ SwiGLU w/ GeLU
Acc (%) 83.8 83.6
Table 9: Reproduced Results of NEPA-B without SwiGLU.
E.6 Additional Quantitative Results
Figure 9 shows additional attention and embedding visualizations for eight ImageNet classes: white wolf (270), giant panda (388), golden retriever (207), balloon (417), volcano (980), loggerhead turtle (33), sulphur-crested cockatoo (89), and space shuttle (812). For each query patch, we display the query map, the attention map over all seen patches, and the cosine-similarity map between the predicted and true embeddings. Across these diverse classes, the model consistently attends to semantically related regions and assigns high similarity to patches on the same object, supporting our claim that NEPA learns object-centric and spatially coherent representations.

### v2

@@
 Generative pretraining has driven major advances in language [radford2018gpt, radford2019gpt2, brown2020gpt3, achiam2023gpt4] and has a long visual lineage.
-Early pixel-level generative pretraining (iGPT) showed transferable features but struggled with very long sequences and weak semantic alignment [chen2020igpt].
+iGPT autoregresses over discrete pixel tokens and learns transferable features, but suffers from long sequences and weak semantic alignment [chen2020igpt].
+AIM series [el2024scalable, fini2025multimodal] extend to the continuous counterpart, supervised on predefined pixel-space targets as continuous patch values.
 Subsequent work introduced discrete tokenizers [van2017vqvae, esser2021vqgan, lee2022rqvae, sun2024llamagen, pang2024randar] to compress images into short code sequences amenable to transformer generative pretraining [esser2021vqgan, ramesh2021dalle, lee2022rqvae], enabling large-scale image generation.
@@
 Modality-agnostic potentials.
-Recent Large Language Models [touvron2023llama] increasingly adopt tied embeddings, where the input and output embedding matrices are shared.
+Recent Large Language Models increasingly adopt tied embeddings, where the input and output embedding matrices are shared.
 This mechanism is effectively equivalent to predicting the next embedding in the latent space, which is precisely the principle underlying our framework.
@@
 We acknowledge the computing resources of Google Cloud Platform and Lambda Labs.
-The authors would like to thank Jiayi Pan and Freda Shi for their valuable feedback.
+We sincerely thank Steven Bucaille for opening a pull request that adds NEPA to the transformers ecosystem.
+We thank Jiayi Pan and Freda Shi for their valuable feedback.
 […]
@@
 In NEPA, the main learned object is the transformer predictor itself: we use a simple shared embedding layer and directly regress the next embedding without negatives or a contrastive head, which makes the architecture easier to scale.
-Relation to GPT [radford2018gpt, chen2020igpt].
-NEPA shares the causal prediction idea of GPT, but works fully in the continuous embedding space. We do not use task-specific decoders, tokenizers, or language heads during pretraining.
+Relation to GPT [radford2018gpt], iGPT [chen2020igpt] and AIM [el2024scalable, fini2025multimodal].
+NEPA shares the causal prediction idea, but works fully in the continuous embedding space.
+We do not use task-specific decoders, tokenizers, or language heads during pretraining.
+iGPT and AIM-style methods supervise against predefined pixel-space targets (discrete pixel tokens or continuous patch values), whereas NEPA supervises in representation space via embedding similarity, a learned, moving target defined by the embedding rather than pixels.
 Relation to JEPA [assran2023jepa].
@@
 All ImageNet experiments use the ILSVRC/imagenet-1k dataset accessed through Hugging Face datasets.
-ADE20K experiments follow the standard recipes and data processing pipelines from mmsegmentation [mmseg2020].
-Our code and model is available at https://sihanxu.github.io/nepa.
+ADE20K experiments follow the standard recipes and data processing pipelines from mmsegmentation [mmseg2020].Our code and model is available at https://github.com/SihanXU/nepa.
 Appendix D Training Details